Convert the speed 1.0000 m/s [meter/second] to mi/h [miles/hour], expressing your answer to 5 significant figures. (Useful info: 1 mile = 5280 feet exactly, and 1 inch = 2.54 cm exactly.) 


1
Expert's answer
2021-08-27T14:39:42-0400
"1\frac{m}{s}=\frac{100}{2.54}\frac{3600}{5280}\frac{1}{12}=2.2369\frac{mi}{h}"